Find all real solutions:
4^x-5(2^x)-24=0

Answer by Alan3354(69443)

4^x-5(2^x)-24=0
-------
2^(2x) - 5*2^x - 24 = 0
--
Sub u for 2^x
u^2 - 5u - 24 = 0
(u-8)*(u+3) = 0
u = -3
2^x = -3 Not allowed
----------
u = 8
2^x = 8
x = 3